40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 940     Entry time: Wed Sep 10 19:53:53 2008
Author: Alberto 
Type: Update 
Category: General 
Subject: abs length experiment 
Update of the last days work on the experiment to measure the absolute length of the cavities.

I'm trying to repeat the same measurement that Koji did on the Y arm, before switching to the X arm.

I switched to the PHD universal box for the PLL control between the main laser and the secondary laser. I found a good gain value for the servo and now I can set the frequency of the beat to any value as long as I do it slowly turning the LO frequency from the knob on the Marconi.

I laid down a 50m BNC cable from the Y end to near the BS chamber, where all the abs length equipment is. I matched the two laser beams changing the alignment of the injection steering mirror at the the dark port on the AP table. I then locked the Y arm cavity. When I first tried to do that, the locking script didn't work because the beam was off of the 'sweet spot' where Rob had set it on Monday. It turned out that aborting the script during one of its previous run, had changed the alignment of the PZT steering mirrors. So with Rob I brought them back near the positions as in the snapshot and then saved a new one with the latest values.

Eventually I could set the beat frequency to the FSR of the arm cavity and saw it in transmission at the ETMY.

Now I'm working on the LabView interface for the GPIB data acquisition board.
ELOG V3.1.3-